# CER-HV: A CER-Based Human-in-the-Loop Framework for Cleaning Datasets Applied to Arabic-Script HTR

### v1

@@ 5.3 Qualitative Examples
Figure 5: Representative examples of flagged samples from the major error categories identified through the CER-HV pipeline: (a) transcription error, (b) segmentation error, (c) orientation error involving a 180° rotation, (d) orientation error involving diagonally oriented text, where a red arrow indicates the intended reading direction, (e) script mismatch, and (f) irrelevant or non-text content. For transcription errors (a), the model prediction is shown to illustrate the mismatch with the ground truth. For structural errors where correction is straightforward (b–c), the corrected image is displayed. Dataset name, image ID, CER, and ground-truth text are shown for all samples. Examples are drawn from multiple datasets (Muharaf, NUST-UHWR, PHTI, and Ajami), highlighting the diverse manifestations of labeling challenges across Arabic-script handwriting corpora.
In addition to noisy samples, the noise detector identified a subset of correctly labeled but visually challenging examples. These “valid but hard” cases often involve text containing numbers or special symbols (e.g., slashes), which are underrepresented in the training data, as well as samples with low visual clarity, unusually long text lines, heavy stroke overlap, or dense diacritics. Although these samples were not corrected or removed, they provide insight into where the model still struggles and help distinguish genuine annotation errors from intrinsic recognition difficulty. Figure 6 illustrates representative cases.
@@ 5.4 Quantitative Impact of Label Errors
Taken together, these results show that even limited amounts of label noise can substantially distort model assessment, emphasizing the need for systematic validation of evaluation sets to ensure reliable and fair benchmarking in HTR. While evaluation-set cleaning is critical for obtaining trustworthy performance estimates, training-set cleaning provides measurable benefits primarily when noise levels are sufficiently high and dominated by structurally disruptive errors that interfere with sequence learning. In settings where training noise is sparse or mild, retraining on cleaned data yields diminishing returns.

### v2

@@ 5.3 Qualitative Examples
Figure 5: Representative examples of flagged samples from the major error categories identified through the CER-HV pipeline: (a) transcription error, (b) segmentation error, (c) orientation error involving a 180° rotation, (d) orientation error involving diagonally oriented text, (red arrow indicates the intended reading direction), (e) script mismatch, and (f) irrelevant or non-text content. For transcription errors (a), the model prediction is shown to illustrate the mismatch with the ground truth. For structural errors where correction is straightforward (b–c), the corrected image is displayed. Dataset name, image ID, CER, and ground-truth text are shown for all samples. Examples are from multiple datasets (Muharaf, NUST-UHWR, PHTI, and Ajami), highlighting the diverse manifestations of labeling challenges across Arabic-script handwriting corpora.
In addition to noisy samples, the noise detector identified a subset of correctly labeled but visually challenging examples. These cases often involve text containing numbers or special symbols (e.g., slashes), which are underrepresented in the training data, as well as samples with low visual clarity, unusually long text lines, heavy stroke overlap, or dense diacritics such as shown in Figure 6. Although these samples were not corrected or removed, they provide insight into where the model still struggles and help distinguish genuine annotation errors from intrinsic recognition difficulty.
@@ 5.4 Quantitative Impact of Label Errors
These results show that even limited amounts of label noise can substantially distort model assessment, emphasizing the need for systematic validation of evaluation sets to ensure reliable and fair benchmarking in HTR. While evaluation-set cleaning is critical for obtaining trustworthy performance estimates, training-set cleaning provides measurable benefits primarily when noise levels are sufficiently high and dominated by structurally disruptive errors that interfere with sequence learning. In settings where training noise is sparse or mild, retraining on cleaned data yields diminishing returns.